OpenStudy (anonymous):

Solve for x: 3(x + 1) = −2(x − 1) + 6 My answer is 4 Am I right?

jimthompson5910 (jim_thompson5910):

No, the answer is not 4

OpenStudy (anonymous):

I think anser is 1

OpenStudy (anonymous):

Oh your right! Thanks! :)

OpenStudy (anonymous):

Hre is how... 3x = -2x+8 5x=8-3 therefor x =1

OpenStudy (anonymous):

welcome :)

OpenStudy (anonymous):

That makes sense! Thank you! :)